# Identifying Information from Observations with Uncertainty and Novelty

### v1

@@ Abstract
A machine learning tasks from observations must encounter and process uncertainty and novelty, especially when it is expected to maintain performance when observing new information and to choose the best fitting hypothesis to the currently observed information. In this context, some key questions arise: what is information, how much information did the observations provide, how much information is required to identify the data-generating process, how many observations remain to get that information, and how does a predictor determine that it has observed novel information? This paper strengthens existing answers to these questions by formalizing the notion of identifiable information that arises from the language used to express the relationship between distinct states. Model identifiability and sample complexity are defined via computation of an indicator function over a set of hypotheses. Their properties and asymptotic statistics are described for data-generating processes ranging from deterministic processes to ergodic stationary stochastic processes. This connects the notion of identifying information in finite steps with asymptotic statistics and PAC-learning. The indicator function’s computation naturally formalizes novel information and its identification from observations with respect to a hypothesis set. We also proved that computable PAC-Bayes learners’ sample complexity distribution is determined by its moments in terms of the the prior probability distribution over a fixed finite hypothesis set.
@@ 1 Introduction
The colloquial definition of “identification” is “to determine an object or concept to which some observations belong.” For example, one could identify a person from their portrait or an artist from their distinctive creative style. The sample complexity of an identification problem is the number of observations required before an identity is determined. Often, we do not make an identification with absolute certainty. In these cases, the sample complexity corresponds to the number of observations required to reach an a-priori level of certainty regarding the identity. Most humans have a powerful memory for faces; identification of a known person from a portrait generally requires only a single image, so we would say that the sample complexity of this problem is one. However, what happens when we can only observe part of the portrait, or when the image is of poor quality? We concern ourselves with this sort of inference problem, where we want to characterize the amount of evidence required to overcome general identification problems. As we will see, an extremely broad class of mathematical and statistical problems can be characterized as identification problems and thus addressed with the formalism we present. This framework will ensure that we have appropriate language and clear assumptions such that the information present in observed phenomena relative to an identification problem can be precisely quantified and discussed.
@@
Thanks to the ubiquity of identification problems, the theory of identification and sample complexity lacks unity and rigor across algorithmic and probabilistic contexts. This becomes extremely clear when we wish to quantify the information that is contained within the model’s description and the observations. Most texts call models ”identifiable” when they may be differentiated based on observations. In the probability literature, this definition frequently includes that their probability distributions differ, as done by van der Vaart (2000, Eq. 5.34). While this definition serves as an appropriate starting point, the nuance in what the observations are or how they affect identifiability is often left implicit or ignored. This has resulted in some recent works that attempts to directly elucidate model identifiability, such as Lewbel (2019). To identify an abstract set of observations with respect to a proposed model is to determine that the observations are consistent with what would be observed were the observations drawn from that model. The theory of computability provides a rigorous foundation for identification in deterministic contexts with full observability, as we can treat this sort of identification problem exactly as we would approach the computation of a boolean indicator function over some domain. We reinforce the connection between computational theory and identifiable models in statistics, along with probably approximately correct identification from statistical learning theory.
@@
Statistical learning theory is primarily concerned with the sample complexity, which is the number of observations required to identify a model from some explicit or implicit set of hypotheses (Shalev-Shwartz and Ben-David, 2014). Sample complexity as a term seems to originate from Haussler (1988) after Valiant (1984) introduced what Angluin (1988) and Angluin and Laird (1988) called and is now popularly known as Probably Approximately Correct (PAC) learning. Valiant (1984) specified PAC with a polynomial computational complexity constraint and an initial sample size constraint to be polynomial in the hypothesis set’s cardinality. Before then, the related term “sample size” that meets the desired empirical risk minimization was considered in the 1966 work and 1971 English translation of Vapnik and Chervonenkis (2015) that founded VC-Theory. The sample size was also included in following complexity analysis works, e.g., Pearl (1978). To better clarify identifiability and sample complexity we formalize these concepts in information-theoretic terms in both algorithmic and probabilistic contexts. This results in a more thorough understanding of the information present in the model’s description and how that information is expressed via the observations. The framework here also allows us to formally address novel information with respect to a model and hypothesis set.
@@
Novelty and the unknown have concerned scholars since antiquity, as evidenced by Plato’s “Allegory of the Cave” (1998, Book VII), Plato’s “Apology” (1999), and the reasoning about the unobserved or “nonapparents” from signs (Asmis, 1984, Ch. 10). There has been a resurgence of academic interest in these topics within the artificial intelligence and machine learning literature over the past three decades, as expressed by Boult et al. (2021); Boult and Scheirer (2024). Practical problems often involve unseen events with respect to the in-sample observations or prior information, and successful machine learning algorithms need to continue to perform well in the presence of such events. Novelty is associated with many subtopics in machine learning, including novelty detection (Markou and Singh, 2003a, b; Pimentel et al., 2014; Faria et al., 2016), open world learning (Bendale and Boult, 2015; Pinto et al., 2022; Doctor et al., 2023; Langley, 2020), and anomaly, outlier, or out-of-distribution detection (Ruff et al., 2021; Chandola et al., 2009). Novelty is also related to the concepts of “discovery” and “emergence” in complex systems as studied in statistical and computational mechanics (Crutchfield, 1994). We refer to “novelty” informally as “something that is unknown to another thing at a moment in time,” which is consistent with prior work and for which we further formalize using set theory and information theory (Shannon, 1948; Cover and Thomas, 1991; MacKay, 2003). Generalization of the predictor’s learned patterns inherently involves incorporation of or robustness to novel information, as evidenced by Zhang et al. (2020). We emphasize the presence of novelty throughout our formalization of identifying information and determine the sample complexity to identify both known and unknown models.
@@ 2 Background, Notation, and Foundation
2.1 Probability: Uncertainty in Variables, Functions, & Processes
We rely upon the common set-theoretic definitions of probability and random variables as by Doob (2012) and Tao (2011). We will often refer to a discrete $(\mathrsfso{X},\Sigma{X})$-valued random variable $X$ with its associated probability triple $(\Omega_{X},\mathit{\Sigma}_{X},P_{X})$ and its sample function $S_{X}:[0,1]\mapsto\Omega_{X}$. Random variables generalize deterministic variables where the probability measures are constant $1$ for the value and 00 everywhere else. We will use them to define the potentially random observed states at each time-step. As we focus on the discrete sample case given computable learning, we define the set of probability distributions over those discrete events. The probability simplex $\mathbb{P}^{k}$ that contains the possible probability mass function, as a vector $p$, of some discrete random variable with support of $k$ mutually exclusive symbols, is defined as
$\mathbb{P}^{k}\triangleq\bigg{\{}p:\sum^{k}_{i=1}p_{i}=1,\leavevmode\nobreak\ p_{i}\in[0,1],\leavevmode\nobreak\ \forall i\in\{1,2,...,k-1,k\}\bigg{\}}$
@@ 3.1 On the Indicator Function and Determining Novelty
We use falsification akin to Popper (2002) as to be able to falsify a known model or hypothesis is crucial to know that the evidence indicates there is an unknown model yet to be discovered that better describes the obseravtions. A useful distinction between undetermined and undefined expressions is that undetermined is yet to be determined, but still can be determined, while undefined is the absence of a description to express an expression’s semantics. Undefined expressions include $\frac{a}{0}:a\in\mathbb{R}$, or the probability of some event given the empirical process upon initialization without any prior occurrences observed. In Section 4, we modify the algorithm that computes an indicator function to measure the distinct states of complex objects represented as strings within some language and extend this concept to asymptotic statistics and PAC learning. Measuring novelty then turns from a Boolean measure to a counting measure and then finally to a general computable measure of information.
@@ 4.1 Definitions
To begin, we provide the definitions of identification and sample complexity used in the rest of the paper. These definitions enable the formalization of information measurements in the output space of a Turing machine $\lambda$ used to determine which parameter $\theta$ in a considered set of parameters $\mathbf{\Theta}$ most probably generated the output $\vec{x\mkern-3.0mu}\mkern 3.0mu^{t}_{1}$, where $t\geq 1$ is the total number of samples and the computation of the sample function is denoted as $\lambda_{S}(\theta,t,Z)=\vec{x\mkern-3.0mu}\mkern 3.0mu^{t}_{1}$, which is covered indepth in Section 4.3. We use $\lambda_{S}(\mathbf{\Theta},t,Z)=\vec{\mathbf{X}\mkern-3.0mu}\mkern 3.0mu^{t}_{1}$ to represent the possible samplings of the different processes within the set $\vec{\mathbf{X}\mkern-3.0mu}\mkern 3.0mu^{t}_{1}$ where each process is described by its corresponding parameter $\theta\in\mathbf{\Theta}$. To simplify our state representation, we consider Turing machines with separate input, memory, and output tapes, which are reducible to single or two tape Turing machines (Minsky, 1967, Ch. 6.3.3). In the following, the input tape will always contain exclusively the description of the output. We primarily focus on finite descriptions and finite memory tapes, but will address how countably infinite descriptions or memory can affect identification. Of particular interest are scenarios with infinite output but finite input, e.g., machines that output computable irrational numbers. We also let the computation time be unbounded so we can focus on when and how a finite sample complexity can be ensured, noting that identification with absolute certainty can only occur when the time to identify is finite, i.e., with finite sample complexity. We start with the common point identifiable parameter of a statistical model as per van der Vaart (2000, Eq. 5.34) and extend that to all the parameters in a set.
@@
Parameters in a set are observationally equivalent with respect to some observed samples if the observed samples do not contain information that identifies the parameters from one another.
@@
Synonyms are “unidentifiable”, “indistinguishable”, and “degenerate”. In physics, two or more states are “degenerate” when they have the same measurement, which is often their energy in quantum mechanics (Sakurai and Napolitano, 2020, Ch. 4.1.3). The measure provides a different enumeration of the distinct states and if two states have the same measurement, then they are observationally equivalent with respect to that measure.
@@
We note that the dimensionality of observations does not appear in these definitions; thus, we can replace the distributions with stochastic processes without loss of generality. Identification requires only the existence of discoverable information that differentiates distinct processes from each other. This information might be contained directly in the observable output. Alternatively, it could be latent, e.g., describing the sampling procedure to generate the observations or indexing a model from a universe of possible models. In either case, we call this the identifying information. We quantify this identifying information by measuring the number of observations that would be required by an effective procedure that determines whether a process is a member of a given set. In the case of latent identifying information as in Section 4.3, we illustrate how that information can be spread over an infinite process and formalize the notion of probabilistic procedure for identification. We can then examine the distribution of the observations required to satisfy a probabilistic decision criterion.
@@
In the context of novelty, what is unknown is determined by exhaustive falsification of what is known. Novelty can then be described in terms of a series of symbol comparisons, where candidates in the known set are falsified one symbol at a time until the location of novel information is encountered.
@@
  • verifies set membership with probability $p$ where $\exists(\vec{X\mkern-3.0mu}\mkern 3.0mu^{*}_{1}\in\vec{\mathbf{X}\mkern-3.0mu}\mkern 3.0mu^{*}_{t},\phi\in\vec{X\mkern-3.0mu}\mkern 3.0mu^{*}_{1}):D(\phi,\vec{x\mkern-3.0mu}\mkern 3.0mu^{t}_{1})\leq\epsilon_{D}$ and $\exists(\vec{X}^{*}_{1}\in\vec{\mathbf{X}\mkern-3.0mu}\mkern 3.0mu^{*}_{1}):2^{t\vec{H\mkern-3.0mu}\mkern 3.0mu(\vec{X\mkern-3.0mu}\mkern 3.0mu)+\log_{2}(p)}\leq P(\vec{X}^{*}_{1}=\vec{x\mkern-3.0mu}\mkern 3.0mu^{t}_{1})\leq 2^{t\vec{H\mkern-3.0mu}\mkern 3.0mu(\vec{X\mkern-3.0mu}\mkern 3.0mu)-\log_{2}(p)}:p\in\mathbb{K},0\leq p\leq 1$,
  • falsifies set membership with probability $q$ where $\forall(\vec{X\mkern-3.0mu}\mkern 3.0mu^{*}_{1}\in\vec{\mathbf{X}\mkern-3.0mu}\mkern 3.0mu^{*}_{t},\phi\in\vec{X\mkern-3.0mu}\mkern 3.0mu^{*}_{1}):D(\phi,\vec{x\mkern-3.0mu}\mkern 3.0mu^{t}_{1})>\epsilon_{D}$ or $\forall(\vec{X}^{*}_{1}\in\vec{\mathbf{X}\mkern-3.0mu}\mkern 3.0mu^{*}_{1}):\big{(}P(\vec{X}^{*}_{1}=\vec{x\mkern-3.0mu}\mkern 3.0mu^{t}_{1})\leq 2^{t\vec{H\mkern-3.0mu}\mkern 3.0mu(\vec{X\mkern-3.0mu}\mkern 3.0mu)+\log_{2}(q)}\big{)}$ or $\big{(}2^{t\vec{H\mkern-3.0mu}\mkern 3.0mu(\vec{X\mkern-3.0mu}\mkern 3.0mu)-\log_{2}(q)}\leq P(\vec{X}^{*}_{1}=\vec{x\mkern-3.0mu}\mkern 3.0mu^{t}_{1})\big{)}:q\in\mathbb{K},0\leq q\leq p$,
@@
An alternative name could be the “identification complexity” as it is fundamentally about how many observations are required in order to identify whether the observations belong to an element in the set; however, we use the common term for consistency with prior works. Also, the set of stochastic processes $\vec{\mathbf{X}\mkern-3.0mu}\mkern 3.0mu^{*}_{1}$ may result in the sample complexity as a distribution over $\mathbb{N}$ as discussed in the following sections and generalizes the deterministic case of a set of strings $\mathrsfso{X}^{*}$.
If $p=0$ or $\epsilon_{D}\rightarrow\infty$, then any model may be used without any observations as certainty and correctness are not of concern, respectively, and accepting $r=1\implies$ accepting $p=0$ and the possibility $\epsilon_{D}\rightarrow\infty$. $p=q=1$ requires absolute certainty. $\epsilon_{D}=0$ requires absolute correctness. Both imply only non-partial identification is accepted. These default values, along with unbounded observations $-\log_{2}(r)\rightarrow\infty$ when $r=0$, may result in a nonhalting program that attempts to compute the sample complexity. We are most interested in when a finite sample complexity occurs without enforcing a maximum with $r>0$, and so we consider the default value $r=0$. Probability $q$ is for probable exhaustive falsification when we cannot have absolute certainty. $p$ and $q$ define two typical sets $\mathcal{A}_{p}\subseteq\mathcal{A}_{q}$, where equality only if $p=q$, for a known process where the most probable sequences all have probability with some deviation from the entropy rate of the process (Cover and Thomas, 1991, Ch. 3). Cover and Thomas (1991, Ch. 16.8) proved that typical sets exist for ergodic stationary stochastic processes. We refer specifically to the sample complexity or its properties with at least probability $p=q=1$, and at most error $\epsilon_{D}=0$ with respect to the Hamming (1950) distance up until asymptotic sample complexity in Theorem 4.3.3.

### v2

@@ Abstract
A machine learning tasks from observations must encounter and process uncertainty and novelty, especially when it is to maintain performance when observing new information and to choose the hypothesis that best fits the current observations. In this context, some key questions arise: what and how much information did the observations provide, how much information is required to identify the data-generating process, how many observations remain to get that information, and how does a predictor determine that it has observed novel information? This paper strengthens existing answers to these questions by formalizing the notion of identifiable information that arises from the language used to express the relationship between distinct states. Model identifiability and sample complexity are defined via computation of an indicator function over a set of hypotheses, bridging algorithmic and probabilistic information. Their properties and asymptotic statistics are described for data-generating processes ranging from deterministic processes to ergodic stationary stochastic processes. This connects the notion of identifying information in finite steps with asymptotic statistics and PAC-learning. The indicator function’s computation naturally formalizes novel information and its identification from observations with respect to a hypothesis set. We also proved that computable PAC-Bayes learners’ sample complexity distribution is determined by its moments in terms of the prior probability distribution over a fixed finite hypothesis set.
@@ 1 Introduction
The colloquial definition of “identification” is “to determine an object or concept to which some observations belong.” For example, one could identify a person from their portrait or an artist from their distinctive creative style. The sample complexity of an identification problem is the number of observations required before an identity is determined. Often, we do not make an identification with absolute certainty. In these cases, the sample complexity corresponds to the number of observations required to reach an a priori level of certainty regarding the identity. Most humans have a powerful memory for faces; identification of a known person from a portrait generally requires only a single image, so we would say that the sample complexity of this problem is one. However, what happens when we can only observe part of the portrait, or when the image is of poor quality? We concern ourselves with this sort of inference problem, where we want to characterize the amount of evidence required to overcome general identification problems. As we will see, an extremely broad class of mathematical and statistical problems can be characterized as identification problems and thus addressed with the formalism we present. This framework will ensure that we have appropriate language and clear assumptions such that the information present in observed phenomena relative to an identification problem can be precisely quantified and discussed.
@@
Thanks to the ubiquity of identification problems, the theory of identification and sample complexity lacks unity and rigor across algorithmic and probabilistic contexts. This becomes extremely clear when we wish to quantify the information that is contained within the model’s description and the observations. Most texts call models “identifiable” when they may be differentiated based on observations. In the probability literature, this definition frequently includes that their probability distributions differ, as done by van der Vaart (2000, Eq. 5.34). While this definition serves as an appropriate starting point, the nuance in how the observations effect model identifiability is not as evident as telling two observations apart and can greatly vary for different hypothesis sets. This complexity has resulted in some recent work that attempts to directly elucidate model identifiability, such as Lewbel (2019). To identify an abstract set of observations with respect to a proposed model is to determine that the observations are consistent with what would be observed were the observations drawn from that model. The theory of computability provides a rigorous foundation for identification in deterministic contexts with full observability, as we can treat this sort of identification problem exactly as we would approach the computation of a Boolean indicator function over some domain. We reinforce the connection between computational theory and identifiable models in statistics by working from computational identification through asymptotic identification up to probably approximately correct identification from statistical learning theory.
@@
Statistical learning theory is primarily concerned with the sample complexity, which is the number of observations required to identify a model from some explicit or implicit set of hypotheses (Shalev-Shwartz and Ben-David, 2014). Sample complexity as a term seems to originate from Haussler (1988) after Valiant (1984) introduced what Angluin (1988) and Angluin and Laird (1988) called and is now popularly known as Probably Approximately Correct (PAC) learning. Valiant (1984) specified PAC with a polynomial computational complexity constraint and an initial sample size constraint to be polynomial in the hypothesis set’s cardinality. Before then, the related term “sample size” that meets the desired empirical risk minimization was considered in the 1966 work and 1971 English translation of Vapnik and Chervonenkis (2015) that founded VC-Theory. The sample size was also included in following complexity analysis works, e.g., Pearl (1978). To better clarify identifiability and sample complexity we formalize these concepts in information-theoretic terms in both algorithmic and probabilistic contexts. This results in a more thorough understanding of the information present in the model’s description and how that information is expressed via the observations.
@@
Novelty and the unknown have concerned scholars since antiquity, as evidenced by Plato’s “Allegory of the Cave” (1998, Book VII), Plato’s “Apology” (1999), and the reasoning about the unobserved or “nonapparents” from signs (Asmis, 1984, Ch. 10). There has been a resurgence of academic interest in these topics within the artificial intelligence and machine learning literature over the past three decades, as expressed by Boult et al. (2021); Boult and Scheirer (2024). Practical problems often involve unseen events with respect to the in-sample observations or prior information, and successful machine learning algorithms need to continue to perform well in the presence of such events. Novelty is associated with many subtopics in machine learning, including novelty detection (Markou and Singh, 2003a, b; Pimentel et al., 2014; Faria et al., 2016), open world learning (Bendale and Boult, 2015; Pinto et al., 2022; Doctor et al., 2023; Langley, 2020), and anomaly, outlier, or out-of-distribution detection (Ruff et al., 2021; Chandola et al., 2009). Novelty is also related to the concepts of “discovery” and “emergence” in complex systems as studied in statistical and computational mechanics (Crutchfield, 1994). We refer to “novelty” informally as “something that is unknown to another thing at a moment in time,” which is consistent with prior work and for which we further formalize using set theory and information theory (Shannon, 1948; Cover and Thomas, 1991; MacKay, 2003). Generalization of the predictor’s learned patterns inherently involves incorporation of or robustness to novel information, as evidenced by Zhang et al. (2020). We emphasize the presence of novelty throughout our formalization of identifying information and determine the sample complexity to identify both known and unknown models. Our theory formally addresses novel information with respect to a model and its hypothesis set, such that a predictor can detect if its hypothesis set is misspecified, which is a problem that plagues Bayesian estimators and PAC-Bayes learning (Kleijn and Vaart, 2006, 2012; Walker, 2013; Wang and Blei, 2019; Lotfi et al., 2022).
@@ 2 Background, Notation, and Foundation
2.1 Probability: Uncertainty in Variables, Functions, & Processes
We model both the data generating process and the predictor as computational stochastic processes. The hypothesis set consists of hypotheses represented as computational models that describe the underlying data generating process of the observations. The predictor is the machine tasked with identifying the best hypothesis in its hypothesis set or exhaustively falsifying each hypothesis given the observations. Parameterized models form a parameterized family of processes typically intended to define distinct models with different output when considering identifiable statistical models, such as families of distributions, but not all models have identifiable parameters by design, which makes determining identification important for learning equivalent hypotheses or their parts. We often refer to identifying a parameter, which is equivalent to identifying a hypothesis, model, or description.
We rely upon the common set-theoretic definitions of probability and random variables as by Doob (2012) and Tao (2011). We will often refer to a discrete $(\mathrsfso{X},\Sigma{X})$-valued random variable $X$ with its associated probability triple $(\Omega_{X},\mathit{\Sigma}_{X},P_{X})$ and its sample function $S_{X}:[0,1]\mapsto\Omega_{X}$. Random variables generalize deterministic variables where the probability measures are constant $1$ for the value and 00 everywhere else. We use them to define the potentially random observed states at each time-step. As we focus on the discrete sample case given computable learning, we define the set of probability distributions over those discrete events. The probability simplex $\mathbb{P}^{k}$ that contains the possible probability mass function, as a vector $p$, of some discrete random variable with support of $k$ mutually exclusive symbols, is defined as
$\vspace{-0.5em}\mathbb{P}^{k}\triangleq\bigg{\{}p:\sum^{k}_{i=1}p_{i}=1,\leavevmode\nobreak\ p_{i}\in[0,1],\leavevmode\nobreak\ \forall i\in\{1,2,...,k-1,k\}\bigg{\}}$
Figure 2: A single time-step of a predictor updating its estimated hypothesis that best describes the observations witnessed up to this point in time.
@@ 3.1 On the Indicator Function and Determining Novelty
If a model has yet to be falsified, then it is not to be discarded as it may still be true, as per the principle of multiple explanations (Asmis, 1984, Ch. 18). We use falsification akin to Popper (2002) as to be able to falsify a known model or hypothesis is crucial to know that the evidence indicates there is an unknown model yet to be discovered that better describes the obseravtions. A useful distinction between undetermined and undefined expressions is that undetermined is yet to be determined, but still can be determined, while undefined is the absence of a description to express an expression’s semantics. Undefined expressions include $\frac{a}{0}:a\in\mathbb{R}$, or the probability of some event given the empirical process upon initialization without any prior occurrences observed. In Section 4, we modify the algorithm that computes an indicator function to measure the distinct states of complex objects represented as strings within some language and extend this concept to asymptotic statistics and PAC learning. Measuring novelty then turns from a Boolean measure to a counting measure and then finally to a general computable measure of information.
@@ 4.1 Definitions
To begin, we provide the definitions of identification and sample complexity used in the rest of the paper. These definitions enable the formalization of information measurements in the output space of a Turing machine $\lambda$ used to determine which hypothesis $\theta$ in a considered set of hypothesis $\mathbf{\Theta}$ most probably generated the output $\vec{x\mkern-3.0mu}\mkern 3.0mu^{t}_{1}$, where $t\geq 1$ is the total number of samples and the computation of the sample function is denoted as $\lambda_{S}(\theta,t,\vec{z\mkern-3.0mu}\mkern 3.0mu)=\vec{x\mkern-3.0mu}\mkern 3.0mu^{t}_{1}$ where $\vec{z\mkern-3.0mu}\mkern 3.0mu\sim Z$ is the results of flipping a fair coin until a sample can be made, which is ignored if deterministic and covered indepth in Section 4.3. We use $\lambda_{S}(\mathbf{\Theta},t,Z)=\vec{\mathbf{X}\mkern-3.0mu}\mkern 3.0mu^{t}_{1}$ to represent the possible samplings of the different processes within the set $\vec{\mathbf{X}\mkern-3.0mu}\mkern 3.0mu^{t}_{1}$ where each process has a description $\theta\in\mathbf{\Theta}$. To simplify our state representation, we consider Turing machines with separate input, memory, and output tapes, which are reducible to single or two tape Turing machines (Minsky, 1967, Ch. 6.3.3). In the following, the input tape will always contain exclusively the description of the output. We focus on finite descriptions and finite memory tapes, but will address how countably infinite descriptions or memory can affect identification. Of particular interest are scenarios with infinite output but finite input, e.g., machines that output computable irrational numbers. We also let the computation time be unbounded so we can focus on when and how a finite sample complexity can be ensured, noting that identification with absolute certainty can only occur when the time to identify is finite, i.e., with finite sample complexity. We start with the common point identifiable parameter of a statistical model as per van der Vaart (2000, Eq. 5.34) and extend that to all the parameters (hypotheses) in a set.
@@
A set of hypotheses are observationally equivalent with respect to some observations if the observations have yet to verify or falsify any hypothesis with respect to some dissimilarity measure $D$ of the observations and hypotheses.
@@
Synonyms are “unidentified”, “undetermined”, “indistinguishable”, and “degenerate”. In physics, two or more states are “degenerate” when they have the same measurement, which is often their energy in quantum mechanics (Sakurai and Napolitano, 2020, Ch. 4.1.3). The measure provides a different enumeration of the distinct states and if two states have the same measurement, then they are observationally equivalent with respect to that measure. In machine learning, we always want to find the observationally equivalent set of best performing models given our performance measures. We can then find the most resource efficient models within that set, and if we can identify correct parts shared across that hypothesis subset, then we can better understand the structure of the underlying concepts.
@@
We note that the dimensionality of observations does not appear in these definitions; thus, we can replace the distributions with stochastic processes without loss of generality. Identification requires only the existence of discoverable information that differentiates distinct processes from each other. This information might be contained directly in the observable output. Alternatively, it could be latent, e.g., describing the sampling procedure to generate the observations or indexing a model from a universe of possible models. In either case, we call this the identifying information. We quantify this identifying information by measuring the number of observations that would be required by an effective procedure that determines whether a process is a member of a given set. In the case of latent identifying information as in Section 4.3, we illustrate how that information can be spread over an infinite process and formalize the notion of a probabilistic procedure for identification. We can then examine the distribution of the observations required to satisfy a probabilistic decision criterion. We now introduce the deterministic case of identification, which we further generalize later.
@@
In the context of novelty, what is unknown is determined by exhaustive falsification of what is known. Novelty can then be described in terms of a series of symbol comparisons, where candidates in the known set are falsified one symbol at a time until novel information is encountered, addressed in Section 4.2.1 after covering the computation of direct identification. Probable novelty is more nuanced and is covered in Section 4.4.2.
For now, we define sample complexity in its most general form for stationary ergodic processes that includes as special cases the deterministic case, asymptotic case, and PAC case of identification, all of which we address in the following subsections.
(a) $p>q\!\leavevmode\nobreak\ \cup\leavevmode\nobreak\ \!|\Theta|>1\!\!\implies\!\!\exists$ undetermined set.
(b) Typical set thresholds over total observations.
Figure 3: The sample complexity typical set thresholds for a single hypothesis of a fair coin with $p=0.7$, $q=0.6$ over $P(\vec{X}^{t}_{1}=\vec{x}^{t}_{1})$. Figure 3(a) depicts the typical and atypical sets at $t=1$ bounding the block entropy in probability space. Figure 3(b) shows the bounds over 10 observations with log scaled probability. Observation sequences within the typical set support the hypothesis. $p$ and $q$ respectively determine the accepted probability of verifying a hypothesis and rejecting all hypotheses. If $p>q$ or $|\Theta|>1$, then there exists an undertermined set of observations where more samples are required to determine set membership.
@@
  • verifies set membership with probability $p$ where $\exists(\vec{X\mkern-3.0mu}\mkern 3.0mu^{*}_{1}\in\vec{\mathbf{X}\mkern-3.0mu}\mkern 3.0mu^{*}_{t},\phi\in\vec{X\mkern-3.0mu}\mkern 3.0mu^{*}_{1}):D(\phi,\vec{x\mkern-3.0mu}\mkern 3.0mu^{t}_{1})\leq\epsilon_{D}$ and $\exists(\vec{X}^{*}_{1}\in\vec{\mathbf{X}\mkern-3.0mu}\mkern 3.0mu^{*}_{1}):2^{t\vec{H\mkern-3.0mu}\mkern 3.0mu(\vec{X\mkern-3.0mu}\mkern 3.0mu)+\log_{2}(p)}\leq P(\vec{X}^{*}_{1}=\vec{x\mkern-3.0mu}\mkern 3.0mu^{t}_{1})\leq 2^{t\vec{H\mkern-3.0mu}\mkern 3.0mu(\vec{X\mkern-3.0mu}\mkern 3.0mu)-\log_{2}(p)},p\in\mathbb{K},0\leq p\leq 1$,
[…]
@@
An alternative name could be the “identification complexity” as it is fundamentally about how many observations are required in order to identify whether the observations belong to an element in the hypothesis set; however, we use the common term for consistency with prior works. Also, the set of stochastic processes $\vec{\mathbf{X}\mkern-3.0mu}\mkern 3.0mu^{*}_{1}$ may result in the sample complexity as a distribution over $\mathbb{N}$ as discussed in the following sections and generalizes the deterministic case of a set of strings $\mathrsfso{X}^{*}$.
If $p=0$ or $\epsilon_{D}\rightarrow\infty$, then any model may be used without any observations as certainty and correctness are not of concern, respectively, and accepting $r=1\implies$ accepting $p=0$ and the possibility $\epsilon_{D}\rightarrow\infty$. $p=q=1$ requires absolute certainty. $\epsilon_{D}=0$ requires absolute correctness. Both imply only non-partial identification is accepted. These default values, along with unbounded observations $-\log_{2}(r)\rightarrow\infty$ when $r=0$, may result in a nonhalting program that attempts to compute the sample complexity. We are most interested in when a finite sample complexity occurs without enforcing a maximum with $r>0$, and so we consider the default value $r=0$. Probability $q$ is for probable exhaustive falsification when we cannot have absolute certainty. Definition 4.1.7 relies on typical sets, which are sets of observation sequences with probability that are representative samples of the block entropy of that probability distribution, i.e., the log probability of the sequences is near that block entropy. $p$ and $q$ define two typical sets $\mathcal{A}_{p}\subseteq\mathcal{A}_{q}$, where equality only if $p=q$, for a known process where the most probable sequences all have probability with some deviation from the entropy rate of the process (Cover and Thomas, 1991, Ch. 3). Figure 3 shows an example of the typical set and atypical sets over the probability measure of the observations. Cover and Thomas (1991, Ch. 16.8) proved that typical sets exist for ergodic stationary stochastic processes. We refer specifically to the sample complexity or its properties with at least probability $p=q=1$, and at most error $\epsilon_{D}=0$ with respect to the Hamming (1950) distance up until asymptotic sample complexity in Theorem 4.3.3, where after we use the relative entropy.